### v1

# Transfer Learning for Spatial Autoregressive Models

Abstract
The spatial autoregressive (SAR) model has been widely applied in various empirical economic studies to characterize the spatial dependence among subjects. However, the precision of estimating the SAR model diminishes when the sample size of the target data is limited. In this paper, we propose a new transfer learning framework for the SAR model to borrow the information from similar source data to improve both estimation and prediction. When the informative source data sets are known, we introduce a two-stage algorithm, including a transferring stage and a debiasing stage, to estimate the unknown parameters and also establish the theoretical convergence rates for the resulting estimators. If we do not know which sources to transfer, a transferable source detection algorithm is proposed to detect informative sources data based on spatial residual bootstrap to retain the necessary spatial dependence. Its detection consistency is also derived. Simulation studies demonstrate that using informative source data, our transfer learning algorithm significantly enhances the performance of the classical two-stage least squares estimator. In the empirical application, we apply our method to the election prediction in swing states in the 2020 U.S. presidential election, utilizing polling data from the 2016 U.S. presidential election along with other demographic and geographical data. The empirical results show that our method outperforms traditional estimation methods.
keywords:
Spatial autoregressive model, transfer learning, U.S. presidential election
journal: Journal of Econometrics
1 Introduction
Spatial dependence among spatial units exists widely in various fields of economics, including urban, real estate, public, agricultural, environmental economics, finance, and so on. The spatial autoregressive (SAR) model (Cliff and Ord,, 1973; Ord,, 1975; Cliff and Ord,, 1975) and its variants are extensively applied in numerous empirical studies concerning spatial competition and spatial spillover effects. For instance, the SAR model is employed to investigate the crime rates in 49 areas of Columbus, Ohio (Anselin,, 1988), budget spillover effects and policy interdependence (Case et al.,, 1993), cigarette demands in various states in the U.S. (Baltagi and Li,, 2004), agricultural land prices in Northern Ireland (Kostov,, 2009), the mutual influence on the stock returns of companies with the same headquarters location (Pirinsky and Wang,, 2006) and the group effect of the dividend policy of listed companies (Grennan,, 2019). Besides, there are also large amounts of applications in other fields where spatial correlations are considered, including political studies (Peters et al.,, 2004), ecology (Ver Hoef et al.,, 2018) and climate sciences (Okunlola et al.,, 2021).
When the sample size of the target data is limited, it is possible to obtain less precise estimates for most statistical models. This issue is particularly relevant in spatial econometrics, where sample sizes can often be quite limited. For instance, in the study investigating crime rates in Columbus, Ohio, the sample size is restricted to just 48 (Anselin,, 1988). This limitation underscores the necessity for additional research aimed at leveraging information from similar datasets to elevate estimation performance. Consequently, there is a pressing need for the exploration of methods to improve the estimation performance of the SAR models through the incorporation of information derived from similar source datasets. Transfer learning is a useful approach to incorporate information from source datasets (Pan and Yang,, 2009; Torrey and Shavlik,, 2010; Zhuang et al.,, 2021; Hu and Zhang,, 2023). The main idea of transfer learning is that the knowledge acquired in one context can be transferred to improve the performance of estimation or prediction when applied to another, even if the two tasks or domains are not identical. In transfer learning, there are typically two main components: (1) The first one is the target domain, to which the knowledge is transferred. The target domain is of our interest but particularly its sample size is limited. (2) The second one is the source domain, from which the knowledge is transferred. In general, the model is pre-trained using the source datasets.
In the recent literature, Bastani, (2021) introduced a two-step estimator that employs high-dimensional techniques to capture biases between the target and source data and has provided an upper bound for the proposed estimators. Li et al., 2022b explored more general scenarios under weaker assumptions and proposed a data-driven method for detecting similarity between the sources and the target. Tian and Feng, (2023) extended the linear model framework to the generalized linear model and introduced an entropy-based method for identifying transferable sources. For more reference on related works, see Cai and Pu, (2019), Cai et al., (2022, 2023), and Li et al., 2022c ; Li et al., 2022a .
In this paper, we introduce a transfer learning framework under the SAR model (tranSAR for short) to enhance the performance of both estimation and prediction. Our study makes three contributions from methodology, statistical theory and empirical application in the literature of the SAR models. First, in methodology, we introduce a two-stage algorithm ($\mathcal{A}$-TranSAR), consisting of a transferring stage and a debiasing stage, to estimate the unknown parameters when the informative source data indexed by $\mathcal{A}$ are known. If we do not know which sources to transfer, a transferable source detection algorithm is proposed to detect informative sources data based on spatial residual bootstrap to retain the necessary spatial dependence. To the best of our knowledge, this is the first work to introduce the transfer learning to the SAR models. Second, in statistical theory, we establish the theoretical convergence rates for the $\mathcal{A}$-TranSAR estimators under the SAR model and theoretically show that the transfer learning can improve the estimation accuracy with the help of the information from the transferable source data. We also derive the detection consistency of the transferable source detection algorithm to avoid the negative transfer when which sources to transfer is unknown. Third, in the empirical application, we apply our method to predict the election outcomes in swing states in the 2020 U.S. presidential election via utilizing polling data from the 2016 U.S. presidential election along with other demographic and geographical data. The empirical results show that our method outperforms traditional spatial econometrics methods.
In the rest of our paper, Section 2 introduces a transfer learning framework for the SAR model and its estimation when the informative source sets are known. When the transferable source data are unknown, a source detection algorithm is also proposed based on spatial residual bootstrap in this section. Section 3 establishes the estimation convergence rates of the transfer learning estimators for the SAR model and derive the detection consistency of the transferable source detection algorithm. In Section 4, we investigate the numerical performance of the proposed methods using Monte Carlo simulations. Section 5 applies the proposed algorithms to the empirical application on U.S. presidential election prediction. We conclude the paper in Section 6.
Notations.
Let $[n]\equiv\{1,\ldots,n\}$. For an index set $\mathcal{A}\subseteq[K]$ and a sequence of positive integers $\{n_{k},k=1,\dots,K\}$, $n_{\mathcal{A}}\equiv\sum_{k\in\mathcal{A}}n_{k}$. For a vector $\mathbf{x}\in\mathbbm{R}^{d}$ and $S$ being a subset of set $[d]$, $\mathbf{x}_{S}$ is the sub-vector of $\mathbf{x}$ with index $S$. The notation $a\lesssim b$ means there exists a positive constant $C\text{ s.t. }a\leq Cb$. The notation $a\vee b$ means $\max(a,b)$, and $a\wedge b$ means $\min(a,b)$. The notation $a_{n}\asymp b_{n}$ means that $a_{n}/b_{n}$ converges to a positive constant. For any random sequences $a_{n}$ and $b_{n}$, $b_{n}=O_{\mathbb{P}}(a_{n})$ means $\forall\epsilon>0,\exists M>0\text{ s.t. }\limsup_{n}\mathbb{P}(|b_{n}|\geq M|a_{n}|)\leq\epsilon$, and $a_{n}\ll b_{n}$ means $|a_{n}|=o_{\mathbb{P}}(|b_{n}|)$, i.e. $\lim_{n\rightarrow\infty}\mathbb{P}(|\frac{a_{n}}{b_{n}}|\geq\varepsilon)=0$ for any $\varepsilon>0$. For a vector $\mathbf{x}={(x_{1},\dots,x_{d})}^{\intercal}\in\mathbbm{R}^{d}$, we define $\|\mathbf{x}\|_{1}=\sum_{j=1}^{d}|x_{j}|$, $\|\mathbf{x}\|_{2}={(\sum_{j=1}^{d}x_{j}^{2})}^{1/2}$, and $\|\mathbf{x}\|_{\infty}=\max_{1\leq j\leq d}|x_{j}|$. For two column vectors $\mathbf{x}$ and $\mathbf{y}\in\mathbbm{R}^{d}$, their inner product is $\langle\mathbf{x},\mathbf{y}\rangle=\mathbf{x}^{\intercal}\mathbf{y}$. For an $n\times n$ matrix $A$, we define $\|A\|_{F}\equiv{(\sum_{i=1}^{m}\sum_{j=1}^{n}\left|a_{ij}\right|^{2})}^{1/2}$, $\|{A}\|_{\infty}\equiv\max_{i}\sum_{j=1}^{n}\left|a_{ij}\right|$, its spectral norm $\|A\|_{2}\equiv\max_{x\in\mathbb{R}^{n}\backslash\{0\}}\frac{\|Ax\|_{2}}{\|x\|_{2}}=\sqrt{\lambda_{\max}(A^{\prime}A)}$, where $\lambda_{\max}(A)$ is the maximum eigenvalue of $A$ and $\lambda_{\min}(A)$ is the minimum eigenvalue of $A$. For an $n\times m$ matrix $A$, $A_{(i,j)}$ represents the element in its $i$-th row and $j$-th column. $A_{(\cdot,j)}$ refers to the $j$-th column of $A$, and $A_{(i,\cdot)}$ represents the $i$-th row of matrix $A$.
2 Transfer Learning for Spatial Autoregressive Model
2.1 Model setting
We consider the spatial autoregressive (SAR) model for the target data,
$Y^{(0)}=\sum_{l=1}^{p}\lambda_{l0}W_{l}^{(0)}Y^{(0)}+\sum_{j=1}^{q}X^{(0)}_{j}\beta_{j0}+V^{(0)},$ (2.1)
where $Y^{(0)}$ is an $n_{0}\times 1$ vector of spatially correlated dependent variables, $W_{l}^{(0)}$ is an $n_{0}\times n_{0}$ non-stochastic spatial weight matrix for each $l\in[p]$, $X_{j}^{(0)}$ is an $n_{0}\times 1$ vector of exogenous regressors independent of the error term $V^{(0)}$ for each $j\in[q]$, $V^{(0)}$ is an $n_{0}\times 1$ vector of independently and identically distributed (i.i.d.) random variables with mean 00 and variance $\sigma^{2}$. Let $\lambda_{0}={(\lambda_{10},\dots,\lambda_{p0})}^{\intercal}$, $\beta_{0}=(\beta_{10},\dots,\beta_{q0})^{\intercal}$ and $\theta_{0}\equiv$ ${({\lambda_{0}}^{\intercal},{\beta_{0}}^{\intercal})}^{\intercal}$.
Similar to general transfer learning settings as in Bastani, (2021)Li et al., (2021)Tian and Feng, (2023), etc., in addition to the target study (2.1), additional samples are obtained from $K$ auxiliary studies (also termed as source datasets), where $K$ is assumed to be fixed. For the $k$-th source dataset, we also consider the SAR model,
$Y^{(k)}=\sum_{l=1}^{p}\lambda_{l0}^{(k)}W_{l}^{(k)}Y^{(k)}+\sum_{j=1}^{q}X^{(k)}_{j}\beta^{(k)}_{j0}+V^{(k)},\ k\in[K],$ (2.2)
where $Z^{(k)}\equiv(Y^{(k)},X_{1}^{(k)},\dots,X_{q}^{(k)})$ is the $k$-th source dataset of sample size $n_{k}$, $W_{l}^{(k)}$ is an $n_{k}\times n_{k}$ non-stochastic spatial weight matrix for each $l\in[p]$, and $V^{(k)}$ is an $n_{k}\times 1$ vector of i.i.d. random variables with mean zero and variance $\sigma^{2}_{k}$, independent of the covariates $X_{j}^{(k)},j\in[q]$. Let $\lambda_{0}^{(k)}={(\lambda_{10}^{(k)},\dots,\lambda_{p0}^{(k)})}^{\intercal}$, $\beta_{0}^{(k)}={(\beta_{10}^{(k)},\dots,\beta_{q0}^{(k)})}^{\intercal}$ and $\omega^{(k)}_{0}\equiv{({\lambda_{0}^{(k)}}^{\intercal},{\beta_{0}^{(k)}}^{\intercal})}^{\intercal}$ for the $k$-th source dataset. Let $\mathbf{X}^{(k)}\equiv(X_{1}^{(k)},\dots,X_{q}^{(k)})$, $\mathbf{W}^{(k)}\equiv({W_{1}}^{(k)},\ldots,{W_{p}}^{(k)})$, and $\mathfrak{D}^{(k)}\equiv(Y^{(k)},\mathbf{X}^{(k)},\mathbf{W}^{(k)})$, $k=0,\dots,K$. In transfer learning, a similar auxiliary model to the target model is called an informative model, and the “similarity” is characterized by the difference between $\omega_{0}^{(k)}$ and $\theta_{0}$, i.e. $\delta^{(k)}\equiv$ $\theta_{0}-\omega^{(k)}_{0}$. Given a tolerance level $h>0$, the index set of the informative auxiliary samples is defined as
$\mathcal{A}\equiv\mathcal{A}_{h}=\{0\leq k\leq K:\|\delta^{(k)}\|_{1}\leq h\},$ (2.3)
where $\|\delta^{(k)}\|_{1}$ is called the transferring level of the $k$-th source dataset. In this context, the value of $h$ is used to balance the bias and the variance introduced by transfer learning. When $h$ is large, more labeled datasets are used, which reduces the prediction variance but increases the bias due to source datasets, and vice versa. We will discuss the value of $h$ in the conditions for theoretical results in Section 3.
2.2 Estimation with known informative source data
In this section, we propose a two-step transfer learning procedure under the SAR model when the index set $\mathcal{A}_{h}$ of the informative auxiliary samples is known, denoted as $\mathcal{A}$-TranSAR. In the first transferring step, we consolidate information from various sources by pooling all available auxiliary data, yielding a preliminary estimator which is typically biased, since $\omega^{(k)}_{0}\neq\theta_{0}$ in general, although they are close. In the second debiasing step, we correct the estimation bias by incorporating the target data via regularization. The details of the two-stage estimation procedure are present in the following.
The transferring stage
The first-step estimator, which is also called pre-training model in the transfer learning literature, with tuning parameter $\lambda_{\omega}$ is defined as
$\widehat{\omega}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\left\{\frac{1}{2n_{\mathcal{A}}}\sum_{k\in\mathcal{A}}\operatorname{\mathit{\ell}}_{1}\left(\mathfrak{D}^{(k)}\mid\omega\right)+\lambda_{\omega}\|\omega\|_{1}\right\},$ (2.4)
where $\operatorname{\mathit{\ell}}_{1}(\cdot\mid\omega)$ is a generic loss function with parameter $\omega$, and $n_{\mathcal{A}}\equiv\sum_{k\in\mathcal{A}}n_{k}$ is the total sample size of all informative source data. For the SAR model, the estimator could be the two-stage least squares (2SLS) estimator (Kelejian and Prucha,, 1998; Lee,, 2003), generalized method of moments estimator (GMM) (Kelejian and Prucha,, 1999; Lee,, 2007; Liu and Lee,, 2010), quasi-maximum likelihood estimator (Lee,, 2004, QMLE) and so on. The main requirement for the first step estimator is a suitable convergence rate discussed in detail in Section 3. Here the tuning parameter $\lambda_{\omega}$ could be zero in low dimensional cases. When the dimension of parameters is greater than the sample size in high dimensional cases, the penalized SAR estimation with nonzero $\lambda_{\omega}$ could produce a sparse estimator, (for example, Higgins and Martellosio, (2023)). See also Zhang and Yu, (2018) for spatial weight matrices selection with model averaging.
Remark 1.
The index set $\mathcal{A}$ may or may not contain the index 00. This depends on the actual application scenarios. In some cases the source data are available, and in some cases, it may be difficult to obtain due to data privacy and security considerations. In the latter case, the owner of the source data can provide the estimation results obtained by the first stage learner without the target data instead of the original source data, and the target task will use this result for the next debiasing stage.
The debiasing stage
In the second step, the bias term estimator $\widehat{\delta}$ with tuning parameter $\lambda_{\delta}$ using the target data is defined as
$\underset{\delta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\widehat{\omega}+\delta\right)+\lambda_{\delta}\|\delta\|_{1}.$ (2.5)
Then, the bias-corrected $\mathcal{A}$-TranSAR estimator is defined as
$\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\widehat{\omega}+\widehat{\delta}.$ (2.6)
The two-stage transfer learning procedure under the SAR model is summarized in Algorithm 1.
Input: target dataset: $\mathfrak{D}^{(0)}$; sources datasets: $(\mathfrak{D}^{(k)},k\in\mathcal{A})$; transferring set: $\mathcal{A}$; and tuning parameters: $\lambda_{\omega}$, and $\lambda_{\delta}$
1
Output: $\hat{\theta}_{\mathcal{A}\textit{-TranSAR}}$
2
3 The transferring stage:
$\widehat{\omega}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\left\{\frac{1}{2n_{\mathcal{A}}}\sum_{k\in\mathcal{A}}\operatorname{\mathit{\ell}}_{1}\left(\mathfrak{D}^{(k)}\mid\omega\right)+\lambda_{\omega}\|\omega\|_{1}\right\}$
4 The debiasing stage:
$\widehat{\delta}=\underset{\delta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\widehat{\omega}+\delta\right)+\lambda_{\delta}\|\delta\|_{1}.$
Output: $\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\widehat{\omega}+\widehat{\delta}$
Algorithm 1 $\mathcal{A}$-TranSAR
Remark 2.
The loss functions $\operatorname{\mathit{\ell}}_{1}$ and $\operatorname{\mathit{\ell}}_{2}$ in the two stages are allowed to be different. In fact, we make no specific requirements on the form of $\operatorname{\mathit{\ell}}_{1}$ as long as the resulting estimator in the transferring stage satisfies the certain convergence rates discussed in Section 3.
In particular, we consider a classical 2SLS procedure in the second stage as an example, which has been widely used for the SAR model to deal with the endogeneity. Given an $n\times d$ ($d\geq p+q$) instrumental variables matrix $Q^{(0)}$ containing $\mathbf{X}^{(0)}$, the first-stage estimator in the 2SLS procedure is
$\widehat{\pi}_{j}=\underset{\pi_{j}\in\mathbb{R}^{d}}{\arg\min}\left\{\frac{1}{2n_{0}}\|{\mathbb{X}}^{(0)}_{j}-Q^{(0)}\pi_{j}\|_{2}^{2}\right\},$ (2.7)
for $j\in[p+q]$, where ${\mathbb{X}}^{(0)}=({W_{1}}^{(0)}Y^{(0)},\dots,{W_{p}}^{(0)}Y^{(0)},\mathbf{X}^{(0)})\in\mathbb{R}^{p+q}$, ${\mathbb{X}}^{(0)}_{j}$ is the $j$-th column of ${\mathbb{X}}^{(0)}$ and $\pi_{j}$’s are coefficients of instrumental variables. Denote the fitted values of ${\mathbb{X}}^{(0)}_{j}$’s by $\widehat{\mathbb{X}}^{(0)}_{j}\equiv Q^{(0)}\hat{\pi}_{j}=P_{Q^{(0)}}{\mathbb{X}}^{(0)}_{j}$, and $\widehat{\mathbb{X}}^{(0)}=(\widehat{\mathbb{X}}^{(0)}_{1},\dots,\widehat{\mathbb{X}}^{(0)}_{p+q})$, where the projection matrix is $P_{Q^{(0)}}\equiv{Q^{(0)}}({Q^{(0)}}^{\intercal}Q^{(0)})^{-1}{Q^{(0)}}^{\intercal}$. In the SAR model, instrumental variables could be $Q^{(0)}=(\mathbf{X}^{(0)},W_{1}^{(0)}\mathbf{X}^{(0)},\dots,W_{p}^{(0)}\mathbf{X}^{(0)})$ with $d=q(1+p)$ columns. We note that if the first column of $\mathbf{X}^{(0)}$ is the intercept term 1, we need to drop it in $W_{j}^{(0)}\mathbf{X}^{(0)}(j\in[p])$ to avoid multicollinearity. See Kelejian and Prucha, (1998) and Lee, (2003) for several alternative choices of instruments. The corresponding loss function for the second stage estimation is
$\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\theta\right)=\|Y^{(0)}-\widehat{\mathbb{X}}^{(0)}\theta\|_{2}^{2}.$ (2.8)
Then, the bias-corrected $\mathcal{A}$-TranSAR estimator can be equivalently obtained by
$\widehat{\theta}_{\mathcal{A}\textit{-TranSAR}}=\underset{\theta\in\mathbb{R}^{p+q}}{\arg\min}\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(\mathfrak{D}^{(0)}\mid\theta\right)+\lambda_{\delta}\|\theta-\widehat{\omega}\|_{1}.$ (2.9)
Remark 3.
For high dimensional settings in the reduced form, the number of instruments is larger than the sample size. We could consider using a penalized 2SLS procedure as
$\widehat{\pi}_{j}=\underset{\pi_{j}\in\mathbb{R}^{d}}{\arg\min}\left\{\frac{1}{2n_{0}}\|{\mathbb{X}}^{(0)}_{j}-Q^{(0)}\pi_{j}\|_{2}^{2}+p_{\lambda_{\pi_{j}}}(\pi_{j})\right\},$ (2.10)
for certain penalty function $p_{\lambda_{\pi_{j}}}(\cdot)$. In the literature, Belloni et al., (2012) proposed the implementation of Lasso and post-Lasso methods to form the first-stage predictions and estimate optimal instruments in a linear instrumental variables model. Fan and Liao, (2014) developed a novel penalized GMM to cope with the endogeneity of a high dimensional model. Cheng and Liao, (2015) studied the problem of choosing valid and relevant moments for GMM estimation and developed asymptotic results for high-dimensional GMM shrinkage estimators that allow for non-smooth sample moments and weakly dependent observations. Zhu, (2018) and Gold et al., (2020) introduced a two-stage Lasso procedure and established some finite-sample convergence rates under different assumptions.
2.3 A transferable source detection algorithm
The $\mathcal{A}$-TranSAR estimator requires the true informative index set $\mathcal{A}$ to be known correctly. However, in most empirical applications, the true informative set $\mathcal{A}$ is unknown. Transferring adversarial auxiliary samples may not improve the performance of the target model and could even cause worse results. The effect of adversarial auxiliary samples is also called negative transfer, meaning that certain inferior results are caused by uninformative source data (Torrey and Shavlik,, 2010; Tian and Feng,, 2023).
An intuitive approach might involve seeking a metric of similarity between each source and the target data. We compute the initial loss based on the target dataset. Subsequently, we combine the target dataset with each source dataset individually to estimate and compute the corresponding losses. If a source dataset is similar to the target dataset, we anticipate two loss values to be comparably close. Hence, we select a critical value to identify which source datasets are informative. For spatial datasets characterized by a network structure, data splitting used in Tian and Feng, (2023) can not be directly applied since it could disrupt this network integrity. As a result, we adopt the spatial residual bootstrap approach as an alternative. This idea is illustrated in details as follows.
First, we compute an initial estimator $\widehat{\theta}_{ini}={({\widehat{\lambda}_{ini}}^{\intercal},{\widehat{\beta}_{ini}}^{\intercal})}^{\intercal}$ using the target samples by certain estimation methods, such as 2SLS, QMLE or GMM, etc. Then, we obtain the residuals $\hat{e}=Y^{(0)}-\mathbb{X}^{(0)}\hat{\theta}_{ini},$ where ${\mathbb{X}}^{(0)}=({W}_{1}^{(0)}{Y}^{(0)},\ldots,{W}_{p}^{(0)}{Y}^{(0)},\mathbf{X}^{(0)})$. Similar to residual bootstrap, we generate three independent copies, $\widehat{e}^{(1)}$, $\widehat{e}^{(2)}$ and $\widehat{e}^{(3)}$ of sample size $n_{0}$ from the empirical distribution of $\widehat{e}$. Then, we generate three-fold bootstrap samples $Y^{(0,r)}$ and $\mathbf{X}^{(0,r)}$ by
$Y^{(0,r)}=$ $\mathbb{X}^{(0)}\widehat{\theta}_{ini}+\widehat{e}^{(r)},~{}~{}\mathbf{X}^{(0,r)}=$ $\mathbf{X}^{(0)},$ (2.11)
for $r=1,2,3$. Let $Z^{(0,r)}=(Y^{(0,r)},\mathbf{X}^{(0,r)})$ for $r=1,2,3$. Let $Z^{(0,-r)}$ be the concatenated bootstrap samples without the $r$the sample, for example, $Z^{(0,-1)}={({Z^{(0,2)}}^{\intercal},{Z^{(0,3)}}^{\intercal})}^{\intercal}$. Define $\mathbf{X}^{(0,-r)},\mathbb{X}^{(0,r)}$ and $\mathbb{X}^{(0,-r)}$ similarly.
Second, for each $r=1,2,3$, we fit the model in (2.1) on the catenated bootstrap samples $Z^{(0,-r)}$ and use the resulting estimator $\widehat{\theta}^{(0,r)}$ to compute the loss value on the remaining bootstrap sample $Z^{(0,r)}$, that is
$\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})\equiv\frac{1}{2n_{0}}\|Y^{(0,r)}-\mathbb{X}^{(0,r)}\widehat{\theta}^{(0,r)}\|_{2}^{2}.$ (2.12)
Then, we compute the average of three loss values, $\mathcal{L}^{(0)}=\sum_{r=1}^{3}\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})/3$, which serves as the baseline loss value for the target dataset. We can also compute the standard deviation of three loss values, denoted by
$\hat{\sigma}=\sqrt{\frac{1}{2}\sum_{r=1}^{3}{(\mathcal{L}^{[r]}(\hat{\theta}^{(0,r)})-\mathcal{L}^{(0)})}^{2}},$ (2.13)
which measures the dispersion of bootstrapped loss values $\mathcal{L}^{[r]}(\widehat{\theta}^{(0,r)})$ relative to their average $\mathcal{L}^{(0)}$. Next, for each $r=1,2,3$, we combine $Z^{(0,-r)}$ with each $k$-th source dataset, fit the model in (2.1) on the newly combined data and compute the loss denoted by $\mathcal{L}^{(k,r)}$ on $Z^{(0,r)}$. Similarly, the average loss for the $k$-th source dataset over three-fold bootstrap samples is $\mathcal{L}^{(k)}$. Consequently, the difference between $\mathcal{L}^{(k)}$ and $\mathcal{L}^{(0)}$ provides a metric of the similarity between the $k$-th source dataset and the target data. Source datasets whose corresponding loss differences below certain threshold are recognized as transferable informative data. Following the idea of Tian and Feng, (2023), we choose $\hat{\sigma}\vee 0.01$ as the detection threshold. Then, the estimated transferable source index set $\widehat{\mathcal{A}}$ is defined as
$\widehat{\mathcal{A}}=\left\{k=1,\ldots,K\mid\left|\mathcal{L}^{(k)}-\mathcal{L}^{(0)}\right|\leq\left(\widehat{\sigma}\vee 0.01\right)\right\}.$ (2.14)
Once the transferable set $\mathcal{A}$ is estimated, we can apply the $\mathcal{A}$-TranSAR in Algorithm 1 to obtain the transfer learning estimator for the SAR model, denoted by $\widehat{\theta}_{\widehat{\mathcal{A}}\textit{-TranSAR}}$, which is termed as the TranSAR estimator. The procedure described above is summarized in Algorithm 2.
1
Input: target: $\mathfrak{D}^{(0)}$; all sources: $(\mathfrak{D}^{(k)},k\in[K])$; tuning parameter: $\lambda_{\theta}^{(k,r)},k\in[K]\cup\{0\},r\in[3]$; an initial estimator, $\widehat{\theta}_{ini}$.
2
Output: $\widehat{\theta}_{\textit{TranSAR}}$ and $\widehat{\mathcal{A}}$.
3
4 Residuals: $\hat{e}=Y^{(0)}-\mathbb{X}^{(0)}\hat{\theta}_{ini}$ ;
5 for $r$ from $1$ to $3$ do
6       Generate an independent copy $\hat{e}^{(r)}$ from $\hat{e}$ and $Y^{(0,r)}$, $\mathbf{X}^{(0,r)}$, $Z^{(0,-r)}$, $\mathbf{X}^{(0,-r)},\mathbb{X}^{(0,r)}$ and $\mathbb{X}^{(0,-r)}$;
7
8      $\widehat{\theta}^{(0,r)}\leftarrow\underset{\theta\in\mathbb{R}^{p+q}}{\arg\min}{\frac{1}{2n_{0}}\operatorname{\mathit{\ell}}_{2}\left(Z^{(0,-r)}\mid\theta\right)+\lambda^{(0,r)}_{\theta}\|\theta\|_{1}}$ ;
9
10      for $k$ from $1$ to $K$ do
11             $\widehat{\theta}^{(k,r)}\leftarrow\underset{\theta\in\mathbb{R}^{p+q}}{\arg\min}\left\{\frac{1}{2\left(n_{k}+2n_{0}\right)}\left\{\operatorname{\mathit{\ell}}_{2}(\mathfrak{D}^{(k)}\cup Z^{(0,-r)}\mid\theta)\right\}+\lambda_{\theta}^{(k,r)}\|\theta\|_{1}\right\}$ ;
12
13       end for
14
15 end for
16
17Calculate the average losses: $\mathcal{L}^{(k)}=\frac{1}{3}\sum_{r=1}^{3}\mathcal{L}^{[r]}(\hat{\theta}^{(k,r)}),k=0,1,\dots,K$ ;
18
19Calculate the standard deviation: $\hat{\sigma}\leftarrow\sqrt{\frac{1}{2}\sum_{r=1}^{3}{(\mathcal{L}^{[r]}(\hat{\theta}^{(0,r)})-\mathcal{L}^{(0)})}^{2}}$ ;
20
21The estimated transferable source index set: $\hat{\mathcal{A}}\leftarrow\left\{k=1,\ldots,K\mid\left|\mathcal{L}^{(k)}-\mathcal{L}^{(0)}\right|\leq\left(\hat{\sigma}\vee 0.01\right)\right\}$;
[…]
23
Algorithm 2 TranSAR
3 Theoretical results
In Section 3, we will establish an estimation convergence rate for the proposed $\mathcal{A}$-TranSAR estimator by assuming $\underline{n}\equiv\min_{k\in[K]\cup\{0\}}n_{k}\rightarrow\infty$. Subsection 3.1 is for general results and Subsection 3.2 is based on the 2SLS estimation method. Finally, we prove the consistency of the transferable source detection algorithm in Subsection 3.3.
For theoretical analysis, we define the population form for the above estimator is defined as
$\omega^{\mathcal{A}}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\sum_{k\in\mathcal{A}}\alpha_{k}\bar{\mathit{\ell}}_{1}^{(k)}(\omega),$ (3.1)
[…]
$\omega^{(k)}=\underset{\omega\in\mathbb{R}^{p+q}}{\arg\min}\ \bar{\mathit{\ell}}_{1}^{(k)}(\omega).$ (3.2)
3.1 Convergence rate when $\mathcal{A}_{h}$ is known
Assumption 1 (Network structure).
The spatial weight matrices $W_{l}^{(k)}$, for $k=0,1,\ldots,K,l=1,\dots,p$, are non-stochastic with zero diagonals, and $p$ is fixed and does not depend on $n_{0}$ or $n_{k}$. $S^{(k)}(\lambda)=I_{n_{k}}-\sum_{l=1}^{p}\lambda_{l}^{(k)}W_{l}^{(k)}$ is invertible for all $\lambda\in\Lambda$, where the parameter space $\Lambda$ is compact. Matrices $W_{l}^{(k)}$ and $S^{(k)}(\lambda)^{-1}$ are uniformly bounded in both row and column sum norms for all $n_{k},k\in[K]\cup\{0\},l\in[p]$ and $\lambda\in\Lambda$. And for all large enough $n_{k}$, $\|W_{l}^{(k)}S^{(k)}(\lambda)^{-1}\|_{F}/\sqrt{n_{k}}$ is bounded away from zero and infinity uniformly over $k\in[K]\cup\{0\},l\in[p]$ and $\lambda\in\Lambda$.
Assumption 2 (Regular conditions for the design matrix and error terms).
(i) Denote $\mathbf{X}^{(0)}\equiv{(\mathbf{x}_{1}^{(0)},\dots\mathbf{x}_{n_{0}}^{(0)})}^{\intercal}$ and the covariance matrix of $\mathbf{x}_{i}^{(0)}$ for each $i$ is $\Sigma_{0}$. For any ${\beta}\in\mathbb{R}^{q},{\mathbf{x}_{i}^{(0)}}^{\intercal}{\beta}\ (i\in[n_{0}])$, are $i.i.d.$ $\kappa_{0}\|{\beta}\|_{2}$ sub-Gaussian with mean zero, and $0<\underline{\psi_{0}}=\lambda_{\min}(\Sigma_{0})\leq\lambda_{\max}(\Sigma_{0})=\overline{\psi_{0}}<\infty$. (ii) The error terms $V_{i}^{(0)}\ (i\in[n_{0}])$ are $i.i.d.$sub-Gaussian with mean zero and variance $\sigma^{2}$. (iii) As $n_{0}$rightarrow +$\infty$+ ∞, $\mathbb{P}\left(\lambda_{\min}(\frac{(\widehat{\mathbb{X}}^{(0)})^{\intercal}\widehat{\mathbb{X}}^{(0)}}{n_{0}})\geq\phi_{0}\right)\rightarrow 1$, where $\phi_{0}$ is a positive constant.
Assumption 3 (Convergence rates of the first step estimator).
The first step estimator $\widehat{\omega}$ minimizing (2.4) satisfies $\|\widehat{\omega}-\omega^{\mathcal{A}}\|_{1}=O_{\mathbb{P}}{(a_{n_{\mathcal{A}}}^{(1)})}$, and $n_{\mathcal{A}}^{-1}\|{{\mathbb{X}}^{\mathcal{A}}}(\widehat{\omega}-\omega^{\mathcal{A}})\|_{2}^{2}=O_{\mathbb{P}}(a_{n_{\mathcal{A}}}^{(2)})$, where $a_{n_{\mathcal{A}}}^{(1)}\rightarrow 0$ and $a_{n_{\mathcal{A}}}^{(2)}\rightarrow 0$, as $n_{\mathcal{A}}\rightarrow\infty$, and ${\mathbb{X}}^{\mathcal{A}}$ is stack of $\mathbb{X}^{(k)}$ for $k\in\mathcal{A}$.
Assumption 4 (Instruments).
When $n_{0}$ is large enough, the IV matrix $Q^{(0)}=\left(\mathbf{X}^{(0)},W_{1}^{(0)}\mathbf{X}^{(0)},\ldots,W_{q}^{(0)}\mathbf{X}^{(0)}\right)$ satisfies $0<\psi_{*}^{(0)}\leq\frac{1}{\sqrt{n_{0}}}\|Q^{(0)}\|_{2}\leq{\psi^{(0)}}^{*}<\infty$ for some constants $\psi_{*}^{(0)}$ and ${\psi^{(0)}}^{*}$. There exist constant coefficient vectors ${\pi}_{j}\in\mathbb{R}^{d}$ for $j\in[p+q]$ such that $\max_{j=1,\dots,p+q}\|\widehat{\pi}_{j}-{\pi}_{j}\|_{2}=O_{\mathbb{P}}(\sqrt{\log(d)/n_{0}})$ and $\|\pi_{j}\|_{2}^{2}$ is uniformly bounded over $j$ and $n_{0}$, where $\widehat{\pi}_{j}\ (j\in[p+q])$ are defined in Eq. (2.7).
Remark 4.
Assumption 1 is standard in spatial econometrics. See Kelejian and Prucha, (1999)Lee, (2004)Lee, (2007) and Lee and Yu, (2014). Meanwhile, we assume the number of spatial weight matrices $p$ is fixed, as either one or very few spatial weight matrices are often used in most empirical studies. For example, in our empirical application, $p=1$. Assumption 2 requires the predictors and error terms to be sub-Gaussian, which is standard in the literature of transfer learning, such as Li et al., 2022b and Tian and Feng, (2023). The eigenvalue condition in Assumption 2 (iii) is widely employed in high-dimensional statistical literature or can be derived from various regular conditions (Mendelson et al.,, 2008; Negahban et al.,, 2009; Raskutti et al.,, 2010; van de Geer and Bühlmann,, 2009; Loh and Wainwright,, 2012). Assumption 3 imposes some conditions on the convergence rate of the first step estimator, which will be verified using a concrete example in Subsection 3.2. Assumption 4 is about the validity of the instrument matrix $Q^{(0)}$ and the refitted value $\widehat{\mathbb{X}}^{(0)}$.
Theorem 1 (Convergence rate of the $\mathcal{A}$-TranSAR estimator).
Let the parameter space be $\Theta=\Lambda\otimes\left\{{\beta}\mid\|{\beta}\|_{2}<M\right\}$ for some positive constant $M$. Under Assumptions 1-4, if the loss function $\operatorname{\mathit{\ell}}_{1}$ is twice differentiable, $\sup_{k\in\mathcal{A}}\|{\Sigma^{\mathcal{A}}}^{-1}\Sigma^{(k)}\|_{1}=\check{C}<\infty$, $s\frac{\log{q}}{n_{0}}=o(1)$, $q\asymp d$, and $\lambda_{\delta}\gtrsim\sqrt{\log{(q)}/n_{0}}$, where
[…]
Denote $S=\operatorname{support}(\theta_{0})$ and $s=|S|$. Then
$\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}\left((\sqrt{\frac{\log{q}}{n_{0}}}h)\wedge(s\frac{\log{q}}{n_{0}})\wedge h^{2}+a_{n_{\mathcal{A}}}^{(2)}\right).$ (3.5)
Remark 5.
When we consider the case of fixed dimension of covariates $q$, under some regularity conditions, $a_{n_{\mathcal{A}}}^{(2)}=\frac{1}{n_{\mathcal{A}}}$. Then, we could obtain $\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}(\min(n_{0}^{-1},h^{2})+n_{\mathcal{A}}^{-1})$, which implies that when the sample size $n_{\mathcal{A}}$ of the informative source data is large enough and $h$ has a suitable convergence rate, the estimation convergence rate of the $\mathcal{A}$-TranSAR estimator is better than that only using the target data.
3.2 Convergence rate under the 2SLS loss
In this section, we verify the requirements in the above theorem for the 2SLS estimation and simplify the convergence rate of the $\mathcal{A}$-TranSAR estimator. As in Eq. (2.8), let $\operatorname{\mathit{\ell}}_{1}(\cdot\mid\omega)$ be the loss function of the 2SLS. To elaborate, let $Y^{\mathcal{A}}$ denote the stack consisting of $Y^{(k)}$ for all $k\in\mathcal{A}$, i.e., $Y^{\mathcal{A}}={({Y^{(k)}}^{\intercal},k\in\mathcal{A})}^{\intercal}$. $\mathbf{X}^{\mathcal{A}}$ and ${\mathbb{X}}^{\mathcal{A}}$ are defined similarly. Then the objective function (2.4) could be rewritten as
$\frac{1}{2n_{\mathcal{A}}}\|Y^{\mathcal{A}}-\widehat{\mathbb{X}}^{\mathcal{A}}\omega\|_{2}^{2}+\lambda_{\omega}\|\omega\|_{1},$ (3.6)
where $\widehat{\mathbb{X}}^{(k)}$ is defined similarly as $\widehat{\mathbb{X}}^{(0)}$ in Subsection 2.2, and $\widehat{\mathbb{X}}^{\mathcal{A}}$ is their stack. Here the IV matrices are $Q^{(k)}$ and $Q^{\mathcal{A}}$. We impose some several similar conditions on $\mathfrak{D}^{(k)}$, $k\in\mathcal{A}$ in the following.
Assumption 5 (Regular conditions for design matrices and error terms on the $k$-th source).
For source $k\in\mathcal{A}$, we have the following assumptions: (i) For the design matrix $\mathbf{X}^{(k)}=(\mathbf{x}_{1}^{(k)},\dots\mathbf{x}_{n_{k}}^{(k)})^{\intercal}$ and any ${\beta}\in\mathbb{R}^{q},{\mathbf{x}_{i}^{(k)}}^{\intercal}{\beta}\ (i\in[n_{k}])$ are $i.i.d.$ $\kappa_{k}\|{\beta}\|_{2}$ sub-Gaussian with mean zero for certain constant $\kappa_{k}$, where $\max_{k\in\mathcal{A}}\kappa_{k}<\infty$. (ii) The error terms $V_{i}^{(k)}\ (i\in[n_{k}])$ are $i.i.d.$sub-Gaussian with mean zero and variance $\sigma_{k}^{2}$. (iii) As $n_{k}\rightarrow+\infty$, $\mathbb{P}\left(\lambda_{\min}\left(\frac{{{\widehat{\mathbb{X}}}^{(k)\intercal}}{\widehat{\mathbb{X}}}^{(k)}}{n_{k}}\right)\geq\phi_{k}\right)\rightarrow 1$, where $\phi_{k}$ is a positive constant.
Assumption 6 (Instruments for the $k$-th source).
For the instruments
$Q^{(k)}=\left(\mathbf{X}^{(k)},W_{1}^{(k)}\mathbf{X}^{(k)},\ldots,W_{q}^{(k)}\mathbf{X}^{(k)}\right),\ k\in\mathcal{A}\cup\{0\},$ (3.7)
when $n_{k}$, the number of rows of $Q^{(k)}$, is large enough, $0<\psi_{*}^{(k)}\leq\frac{1}{\sqrt{n_{k}}}\|Q^{(k)}\|_{2}\leq{\psi^{(k)}}^{*}<\infty$ for some constants $\psi_{*}^{(k)}$ and ${\psi^{(k)}}^{*}$. There exist constant coefficient vectors $\pi_{j}^{(k)}\in\mathbb{R}^{d}$ for $j\in[p+q]$ such that $\max_{j=1,\dots,p+q}\|\hat{\pi}_{j}-\pi_{j}^{(k)}\|_{2}=O_{\mathbb{P}}(\sqrt{\log(d)/n_{k}})$ and $\|\pi_{j}^{(k)}\|_{2}^{2}$ is uniformly bounded over $j$ and $k$.
Under Assumptions 5 and 6, the convergence rate of the $\mathcal{A}$-TranSAR estimator is summarized in the following Theorem 2.
Theorem 2.
If Assumptions 1245, and 6 hold, $\sup_{n_{k}}{\mathbf{B}^{(k)}}<\infty$ with ${\mathbf{B}^{(k)}}\equiv 1\vee\|\beta_{0}^{(k)}\|_{2}$, $K$ is a fixed, $q\asymp d$, $s\frac{\log{q}}{n_{0}}=o(1)$, $s^{2}\frac{\log{n_{0}}}{n_{0}}=o(1)$, $n_{\mathcal{A}}>n_{0}$, $q^{2}h=O(1)$, $\lambda_{\delta}\gtrsim\sqrt{(\log{q})/n_{0}}$ and $\lambda_{\omega}\gtrsim\sqrt{(\log{q})/n_{\mathcal{A}}}$, where $S=\operatorname{support}(\theta_{0})$ and $s=|S|$, then we have
$\|\widehat{\theta}-\theta_{0}\|_{2}^{2}=O_{\mathbb{P}}\left((\sqrt{\frac{\log{q}}{n_{0}}}h)\wedge({\frac{s\log{q}}{n_{0}}})\wedge h^{2}+(\frac{\log{n_{A}}}{n_{\mathcal{A}}}h^{2})\wedge\frac{\log^{3}{n_{A}}}{n_{\mathcal{A}}^{3}}+\frac{s\log{q}}{n_{\mathcal{A}}}\right).$ (3.8)
Theorem 2 indicates that if $h\ll s\sqrt{\log q/n_{0}}$ and $n_{\mathcal{A}}\gg n_{0}$, the $\mathcal{A}$-TranSAR estimator is better than the classical 2SLS estimator based on only the target data, which demonstrates the usefulness of transfer learning in the SAR models.
3.3 Detection consistency
Finally, we will show that our proposed transferable source detection algorithm is consistent even for a high dimensional model, i.e., $q\gg n_{0}$.
Assumption 7 (Identifiability condition).
The tolerance level $h=o(1)$ as $\underline{n}\equiv\min_{k\in[K]\cup\{0\}}n_{k}\rightarrow\infty$. For any $k\in[K]\cup\{0\}$ and $r\in[3]$, $q\log(q)\|\widetilde{\theta}^{(k)}-\widehat{\theta}^{(k,r)}\|_{1}=o_{\mathbb{P}}(1)$. For $k\in\mathcal{A}^{c}\equiv[K]-\mathcal{A}$, we control the signal power as
$\underline{\lambda}\|\theta_{0}-\widetilde{\theta}^{(k)}\|_{2}^{2}\geq$ $\|{\theta}_{0}-\widetilde{\theta}^{(k)}\|_{1}+\bar{C}\left[(\|\theta_{0}-\widehat{\theta}^{(0,r)}\|_{1})\vee 1\right]-$
$\frac{\sigma^{2}}{2n_{0}}\operatorname{tr}\{[(S(\tilde{\lambda}^{(k)})-S)S^{-1}]^{\intercal}[(S(\tilde{\lambda}^{(k)})+S)S^{-1}]\},$ (3.9)
and the gap between $\theta_{0}$ and $\widetilde{\theta}^{(k)}$,
$\varphi_{k}\equiv\sqrt{\frac{1}{n_{0}}}(\|\theta_{0}-\widetilde{\theta}^{(k)}\|_{1}+1)^{2}=o(1),$ (3.10)
where $\bar{C}>0$ is a constant and $\underline{\lambda}\equiv\liminf_{n_{0}\rightarrow\infty}\lambda_{\min}(\frac{1}{2n_{0}}\mathbb{E}{{\mathbf{X}}^{\star}}^{\intercal}{\mathbf{X}}^{\star})>0$, with ${\mathbf{X}}^{\star}\equiv(\mathbf{W}^{(0)}(\mathbf{I}_{p}\otimes(S^{-1}{\mathbf{X}}^{(0)}{\beta}_{0})),{\mathbf{X}}^{(0)})$.
Remark 6.
Conditions (7) and (3.10) guarantee a suitable gap between the population-level coefficients in set $\mathcal{A}^{c}$ and the true coefficient of the target model. Intuitively, assuming an appropriate signal strength is necessary to ensure the consistency of estimators. On the one hand, a very weak signal makes it challenging for the algorithm to differentiate informative and uninformative sources. On the other hand, an excessively strong signal can lead the first step estimator to mimic the model structure of the source, thereby hindering accurate identification.
Next, we establish the detection consistency of $\widehat{\mathcal{A}}$ in the following theorem.
Theorem 3 (Detection consistency of $\widehat{\mathcal{A}}$).
If Assumptions (1)-(4) and (7) are satisfied, $\|\beta_{0}\|_{2}<\infty$ and $\frac{\log q}{n_{0}}=o(1)$, then the automatically detected transferable index set $\widehat{\mathcal{A}}$ in Eq. (2.14) is consistent, i.e.,
[…]
4 Simulation
In this section, we design several simulation experiments to investigate the performance of the proposed method (denoted as TranSAR). We compare the finite sample performance of our estimator with the classical 2SLS estimation (which is fitted only using the target dataset, denoted as SAR), the ${\mathcal{A}}$-TranSAR estimator (which is fitted based on the true informative index set $\mathcal{A}$, denoted as Oracle TranSAR), pooled-estimator (which is fitted using all the source datasets as the transferable sets, denoted as $[K]$-TranSAR), and the generalized linear model transfer learning algorithm (denoted as glmtrans) proposed by Tian and Feng, (2023). The glmtrans ignores the spatial information and can be considered as a baseline estimator to show that it is necessarily important to consider the spatial correlations if they exist.
We generate the data according to Models (2.1) and (2.2). In these models, $n_{0}=256,n_{1}=n_{2}=\cdots=n_{K}=100,p=1,q=200,K=20$. The covariates $\mathbf{x}_{i}^{(k)}$ are generated by multivariate normal distribution with covariance matrix $\Sigma\in\mathbb{R}^{q\times q}$, whose the $(j,j^{\prime})$-th entry is denoted as $\Sigma_{j,j^{\prime}}$ for all $0\leq k\leq K$. We consider three designs for $\Sigma$: (1) $\Sigma_{j,j^{\prime}}=1$ for $j=j^{\prime}$ and $\Sigma_{j,j^{\prime}}=0$ otherwise, (2) $\Sigma_{j,j^{\prime}}=0.5^{|j-j^{\prime}|}$, (3) $\Sigma_{j,j^{\prime}}=0.9^{|j-j^{\prime}|}$. We consider two different distributions for the i.i.d. error terms: (1) the standard normal distribution $N(0,1)$, and (2) the $t_{2}$ distribution. The spatial units are assumed to be located on a square grid, and we set up several candidate spatial weight matrices as follows: (1) the first $W_{1}$ is the spatial weight matrix where spatial units interact with their neighbors on the left and right sides; (2) the second $W_{2}$ is the spatial weight matrix where spatial units interact with their neighbor above and below; (3) the rest of the spatial weight matrices are assumed to be matrices where spatial units interact with their second-nearest, third-nearest, and so on, denoted as $W_{3},\ldots,W_{N}$, where $N<\sqrt{n_{0}}$. In each dataset, we randomly draw spatial weight matrices from these candidate matrices. The weight matrices ${\{W_{k,l}\}}_{k\in\{0\}\cup[K],l\in[p]}$ are drawn randomly from $\{W_{1},\ldots,W_{N}\}$ without replacement. In Models (2.1) and (2.2), we consider coefficients settings as follows. In the target Model (2.1), $\lambda_{0}$ $=0.4$ and $\beta_{0}$ $=(\mathbf{1}_{3}^{\intercal},\mathbf{0}_{q-3}^{\intercal})^{\intercal}\in\mathbb{R}^{q}$. In Model (2.2),
$\lambda_{0}^{(k)}=\lambda_{0}\mathbbm{1}\{k\in\mathcal{A}\}-\lambda_{0}\mathbbm{1}\{k\in\mathcal{A}^{c}\},$
$\beta_{j0}^{(k)}=\beta_{j0}-0.05\mathbbm{1}\{j\in\mathcal{H}_{k},k\in\mathcal{A}\}-2\mathbbm{1}\{j\in\mathcal{H}_{k},k\in\mathcal{A}^{c}\},j=1,\dots,q,$
where $\mathcal{H}_{k}$ is a random subset of $[q]$ with $|\mathcal{H}_{k}|=H$ if $k\in\mathcal{A}$ (i.e., $H$ is the number of different coefficients in the informative set), or $|\mathcal{H}_{k}|=\frac{q}{2}$ if $k\notin\mathcal{A}$. So, on the informative datasets, the differences of different $\beta_{j0}^{(k)}$’s are 0.05; on uninformative datasets, the ones of different $\beta_{j0}^{(k)}$’s are 2.
We compute the root mean squared errors (RMSE),
$\mbox{RMSE}=\sqrt{\frac{1}{R}\sum_{r=1}^{R}\|{\theta_{0}}-{\widehat{\theta}}^{[r]}\|^{2}},$ (4.1)
where ${\widehat{\theta}}^{[r]}$ is the estimator of the $r$-th replication, based on the 2SLS loss with $q=200$ and display them under different settings in Figure 1. In order to demonstrate the advantages of transfer learning, we let the size of informative auxiliary datasets $|\mathcal{A}|$ increase from 00 to $20$ to show the decreasing trend in RMSE. As the number of informative sets increases, compared to the classical 2SLS estimation and the glmtrans method, we can see that $\widehat{\mathcal{A}}$-TranSAR and ${\mathcal{A}}$-TranSAR can significantly reduce the RMSE, resulting in more precise estimation. Additionally, we can see that the $\widehat{\mathcal{A}}$-TranSAR performs almost the same as the oracle transfer learning method $\mathcal{A}$-TranSAR. That means, our proposed detection method is able to select informative sets with a high probability. Moreover, from Figure 1, when the number of informative sets is small, the negative transfer effect of the pooled method, $[K]$-TranSAR becomes apparent. Hence, it is necessary to select informative sets accurately. The same conclusion applies to all three types of design matrices and the three different settings for $H$. Detailed and comprehensive simulation results are compiled in Table 1 for covariates designs type 1.
Figure 1: RMSE of Three Types of Covariates Designs with Normal Error Distribution, and $H=0,5,10$
Note: The size of informative auxiliary datasets $|\mathcal{A}|$ increases from 00 to $20$.
Table 1: The RMSE of Various Estimators with Standard Normal Distribution Error Terms
H=0 H=5 H=10
$|\mathcal{A}_{h}|$ Method RMSE of $\widehat{\lambda}$ Total RMSE RMSE of $\widehat{\lambda}$ Total RMSE RMSE of $\widehat{\lambda}$ Total RMSE
Covariates design 1
$|\mathcal{A}_{h}|$ = 0 SAR 0.01931 0.49616 0.01593 0.44895 0.01507 0.43567
$[K]$-TranSAR 0.03309 0.51944 0.03755 0.69910 0.03892 0.70725
glmtrans NA 0.66184 NA 0.63342 NA 0.63342
TranSAR 0.02088 0.51506 0.01574 0.45135 0.01574 0.44359
Oracle TranSAR 0.01938 0.49908 0.01556 0.44738 0.01638 0.45203
$|\mathcal{A}_{h}|$ = 10 SAR 0.01536 0.43773 0.01568 0.44799 0.01633 0.45635
$[K]$-TranSAR 0.07954 0.32422 0.07143 0.33693 0.06139 0.43960
glmtrans NA 0.59838 NA 0.60811 NA 0.58176
TranSAR 0.00868 0.07301 0.00842 0.07117 0.00842 0.07073
Oracle TranSAR 0.00873 0.07296 0.00835 0.07138 0.00844 0.07052
$|\mathcal{A}_{h}|$ = 20 SAR 0.01552 0.44279 0.01598 0.44674 0.01568 0.44669
$[K]$-TranSAR 0.00643 0.04898 0.00692 0.04951 0.00614 0.05267
glmtrans NA 0.59203 NA 0.58152 NA 0.59955
TranSAR 0.00648 0.04909 0.00693 0.04963 0.00613 0.05275
Oracle TranSAR 0.00642 0.04853 0.00698 0.05011 0.00614 0.05288
Covariates design 2
$|\mathcal{A}_{h}|$ = 0 SAR 0.01284 0.34807 0.01256 0.34623 0.01203 0.33645
$[K]$-TranSAR 0.03664 0.52357 0.03667 0.53178 0.03734 0.52947
glmtrans NA 0.73390 NA 0.73390 NA 0.73390
TranSAR 0.01305 0.35195 0.01173 0.33473 0.01309 0.35113
Oracle TranSAR 0.01188 0.34144 0.01279 0.34645 0.01280 0.34603
$|\mathcal{A}_{h}|$ = 10 SAR 0.01254 0.34102 0.01179 0.34173 0.01204 0.33671
$[K]$-TranSAR 0.02236 0.38791 0.02557 0.27970 0.03508 0.52258
glmtrans NA 0.65766 NA 0.74894 NA 0.66023
TranSAR 0.00615 0.08292 0.00586 0.07913 0.00561 0.07733
Oracle TranSAR 0.00605 0.08292 0.00581 0.07938 0.00567 0.07754
$|\mathcal{A}_{h}|$ = 20 SAR 0.01314 0.35434 0.01307 0.35456 0.01282 0.34617
$[K]$-TranSAR 0.00465 0.05552 0.00479 0.05768 0.00418 0.05992
glmtrans NA 0.67016 NA 0.68051 NA 0.66777
TranSAR 0.00462 0.05539 0.00479 0.05770 0.00420 0.06011
Oracle TranSAR 0.00462 0.05542 0.00479 0.05763 0.00418 0.06007
Covariates design 3
$|\mathcal{A}_{h}|$ = 0 SAR 0.01044 0.44594 0.01052 0.44397 0.01006 0.43836
$[K]$-TranSAR 0.01525 1.07038 0.01479 1.07295 0.01464 1.08117
glmtrans NA 1.02237 NA 1.02237 NA 1.02237
TranSAR 0.01000 0.44203 0.00959 0.43640 0.01076 0.45431
Oracle TranSAR 0.01002 0.43889 0.01011 0.44119 0.01053 0.44673
$|\mathcal{A}_{h}|$ = 10 SAR 0.01070 0.44301 0.01078 0.44921 0.00960 0.43615
$[K]$-TranSAR 0.02097 1.84294 0.02474 0.82913 0.02935 1.75217
glmtrans NA 0.88731 NA 1.23627 NA 0.94350
TranSAR 0.00508 0.17931 0.00512 0.17545 0.00493 0.15996
Oracle TranSAR 0.00508 0.17984 0.00507 0.17543 0.00493 0.15873
$|\mathcal{A}_{h}|$ = 20 SAR 0.01058 0.44641 0.01009 0.43955 0.00990 0.43987
$[K]$-TranSAR 0.00381 0.11680 0.00413 0.12832 0.00340 0.12749
glmtrans NA 0.81752 NA 0.95240 NA 0.85179
TranSAR 0.00383 0.11714 0.00413 0.12769 0.00340 0.12760
Oracle TranSAR 0.00382 0.11678 0.00413 0.12836 0.00341 0.12748
Notes: The RMSE of $\widehat{\lambda}$ is the RMSE of the estimated coefficients of the $\lambda_{0}$. The total RMSE is the RMSE of all the estimated coefficients of the covariates. “NA” indicates that “glmtrans” is not applicable for the estimation of $\lambda_{0}$.
5 Empirical Study
In recent years, amounts of Geographic Information Science (GIS) data have been used in election campaigns (Peters et al.,, 2004). Geography is the common denominator of where voters live, how residents engage in the voting process, how officials manage elections, and where elections campaign runs. Based on population data and geographic information, election district boundaries are resettled to draft multiple proposed plans and demonstrate the effects of several approaches, for instance, moving a border to divide a community of interest. Political campaigns make full use of geographic information and related tools to obtain underlying support from voters, especially the ones from swing states.
Shaping voter preference in swing states is an essential problem, especially for the prediction target. Including spatial information in data analysis brings a novel insight into the prediction. In our empirical study, we focus on the 2020 U.S. presidential election results of each county for swing states by incorporating the spatial relationship among counties. We try to predict the election result using the polls from the 2016 election U.S. presidential election and some other demographic data with geographic information. It is worth noting that swing states have undergone historical shifts over time. The process of identifying swing states in previous elections typically involves an assessment of the closeness of the vote margins in each state. This analysis incorporates historical election outcomes, opinion polls, political trends, recent developments since the previous election, and the specific attributes, strengths, or vulnerabilities of the candidates in contention. We select six swing states for consideration from mainstream journals and news media, including Florida, Georgia, Michigan, Minnesota, North Carolina, and Ohio.
The digital boundary definitions of the United States congressional district are available at https://cdmaps.polisci.ucla.edu/, and the cartographic boundary shape files of counties in the USA are available at the United States Census Bureau (https://www.census.gov/geographies/mapping-files/time-series/geo/cartographic-boundary.2020.html). The spatial weight matrix is constructed using above two geographic datasets. We apply the “queen” contiguity-based spatial weights and row-normalized. The “queen” criterion defines neighbors as spatial units sharing a common edge or a common vertex. The response is the difference of support at the county level for presidential candidates from both parties in 2016, available at https://github.com/tonmcg/US_County_Level_Election_Results_08-20. The county-level demographic information forms the predictors, collected at https://www.kaggle.com/datasets/benhamner/2016-us-election. Among all states, we exclude Alaska, Hawaii, and Washington, D.C.. After regular preprocessing, there are 3106 counties and 45 county-level predictors, among 48 U.S. states. Then we add the pairwise interaction terms between predictors as new predictors. Finally, there are 1035 county-level predictors.
Each swing state is regarded as a target, and all the other states are considered as data sources. We run the proposed estimation method, TranSAR, using the polls from the 2016 U.S. presidential election and some other demographic data with geographic information. We also estimate the SAR model with Lasso penalty (due to high dimensional covariates) using only the target datasets (denoted as SAR) to compare the empirical performance. Due to the inherent stochastic nature of the resampling procedure for the algorithm, we conduct 20 replications for both methods. The average prediction RMSE of the county-level vote results of the county-level 2020 U.S. presidential election is compared at Table 2. For all swing states, the proposed TranSAR method performs better than the traditional SAR in terms of RMSE.
Table 2: The Average RMSE of the Predicted County-level Rate of Support in the 2020 Presidential Election.
Florida Georgia Michigan Minnesota North Carolina Ohio
SAR 0.10723 0.10995 0.10494 0.19211 0.10118 0.10984
tranSAR 0.09417 0.09064 0.08033 0.09646 0.10966 0.08297
Moreover, we also consider the practical rules of the U.S. presidential election, which operate on a “winner-takes-all” basis. In this context, the winner party secures all the state’s electoral votes. Given the “winner-take-all” framework of the U.S. presidential election, we focus on predicting the winner party in each state. Leveraging the county-wise population, we compute the final difference of support in the state-level endorsement for presidential contenders belonging to the respective parties. We calculate the rate of support for state-level elections using the following formula,
[…]
Then, we compare the predicted rate of state level support and the true one in $2020$ election, showing in Table 3. It demonstrates that the transfer learning under the SAR framework is able to effectively improve the state-level prediction.
Table 3: Predicted Bias of the Predicted State-level 2020 Election Results.
Florida Georgia Michigan Minnesota North Carolina Ohio
SAR -0.01989 -0.08110 -0.07157 -0.04189 -0.01782 -0.05765
tranSAR 0.00546 -0.04186 0.03522 -0.02228 -0.02163 -0.03189
The calculated ratios of correct predictions out of the twenty election result forecasts are collected in Table 4. As shown in Table 4, the method we proposed exhibits better performance, especially in predicting election outcomes in Michigan and Minnesota. Simultaneously, it is noteworthy that for the election predictions in Georgia, neither of the methods works well. It can be explained by the historical context: from $1972$ to $2016$, Georgia consistently supported the Republican Party, except for Democratic candidates from the southern region. However, the state has become an increasingly competitive battleground. In 2020, the Democratic candidate, Joe Biden, secured a victory over Donald Trump by a narrow margin of $0.2\%$. This marks a significant departure from the history, as the Democrats secured the presidential election with a slim advantage. Evidently, Georgia has become an outlier in this regard. It is evident that predicting the ultimate election outcome for a single state is significantly more challenging than forecasting the rate of state-level support. Nevertheless, we have made commendable progress in this regard. To provide a comprehensive demonstration of the superiority of the proposed method, we treat each state as the target once a time. Subsequently, we conduct both the tranSAR and SAR processes, and the outcomes have been summarized into Table LABEL:tab:_emp_compare_ele_result_for_all_states for reference, which can be found in LABEL:sec:_addtional_empirical_results. In conclusion, our tranSAR method consistently outperforms the traditional SAR method in the majority of cases, yielding superior prediction results.
Table 4: The Ratios of Correct Predictions for Election Result of 2020 Election Using the 2016 Election Data as Training Data.
Florida Georgia Michigan Minnesota North Carolina Ohio
SAR 100% 0% 35% 90% 100% 100%
tranSAR 100% 0% 100% 100% 100% 100%
In addition, we attempt to visualize transferable states, conducting a statistical analysis of the results after $20$ repetitions. For a swing state, we record the frequency of occurrences for all transferable states and depict states with more than $10$ occurrences (50%) on the map in green color, see Figure 2.
Figure 2: The Detected Transferable States for Each Swing State.
As observed in the figure, some transferable states exhibit geographical information signals. However, for some targets, the spatial or geographical correlations may not be readily apparent. This is a commonplace occurrence. With the development of the economy, culture, and scientific technology, regional interconnections are no longer confined solely to spatial proximity. Latent economic and political affiliations become pivotal factors. We have also created a total visualization map, which is provided in Figure LABEL:fig_emp_transferable_states_map_top_50_all_states at LABEL:sec:_addtional_empirical_results for reference. In this study, we treat each state in the United States as the target once a time, and employ transfer learning techniques similar to those described earlier to construct these maps.
6 Conclusion
We proposed a new transfer learning method under the SAR models, aiming to address the challenges of small sample sizes and spatial dependencies. We derived the theoretical convergence rates for a two-step $\mathcal{A}$-TranSAR estimator when the informative source data sets are known, which is more efficient than the traditional SAR estimator as long as the sample size of the informative source data sets is large enough. Then, we introduced an algorithm for detecting the informative source datasets based on spatial residual bootstrap and proved its detection consistency. The empirical application of this methodology in the prediction of U.S. election results in swing states utilizing demographic and geographical data demonstrates its potential to improve the traditional spatial methodology. Further research might focus on tests for the detection of informative sets and improving the accuracy of coefficients estimation within the transfer learning framework.
Acknowledgements
This work is supported by National Key R&D Program of China (2022YFA1003800), the National Natural Science Foundation of China (NNSFC) (71988101, 72073110, 72333001, 12231011), National Statistical Science Research Grants of China (Major Program 2022LD08), and the 111 Project (B13028).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT 4.0 in order to polish the writing. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.